OpenStudy (thecalchater):

Calculus 5 Questions Walk through 4 check 1?

OpenStudy (thecalchater):

1. B 2. 3. 4. C? 5.

OpenStudy (kainui):

I'll help with questions about concepts or ideas you have doubts on

OpenStudy (thecalchater):

I am not sure on 2 3 or 5 i am pretty sure i did 1 and 4 right. Can you help? and walk me through 2 3 and 5?

OpenStudy (thecalchater):

@mathstudent55 @robtobey @agent0smith @welshfella

OpenStudy (thecalchater):

@IrishBoy123

OpenStudy (zzr0ck3r):

Im at a coffee shop( about to leave) and this would take to long. Ill be home in an hour and will check back

OpenStudy (thecalchater):

ok thanks!

OpenStudy (welshfella):

question 4 looks good for C.

OpenStudy (irishboy123):

it's C for the one that goes \(y' = x(6+y)\). so you separate it, etc etc. so but that is pretty sh1tty, just dumping all that crap here. must do better !!!!!!!

OpenStudy (thecalchater):

lol so 1 and 4 look good. Can you guys walk me through 2?

OpenStudy (thecalchater):

@Zarkon

OpenStudy (thecalchater):

@ganeshie8

OpenStudy (welshfella):

number 5 is exponential decay half life is 1599 so we can write M = 0.5M e^(-kt) M = 0.5M e^(-1599*k) from this you can calculate the values of k then use the formula with t = 15000

OpenStudy (welshfella):

- and 11 for the mass

OpenStudy (thecalchater):

so M=.5M e^(15000*k)

OpenStudy (welshfella):

no i got the masses wrong sorry mass = half the original mass after 1599 years so its 0.5M = M e^(-1599k)

OpenStudy (welshfella):

agree?

OpenStudy (thecalchater):

ok but we need to find 15000 years so take that and plug it into where -1599 is and then what?

OpenStudy (thecalchater):

but it would be -15000 right?

OpenStudy (welshfella):

no you solve for k first using -1599 which is the half life then use the value of k to find mass after 15000 years

OpenStudy (thecalchater):

Well using 11 as the mass like it says i got -.00043348

OpenStudy (thecalchater):

@Zarkon

OpenStudy (welshfella):

k does work out to that so required mass after 15,000 = 11 * e^(-0.00043349* 15,000)

OpenStudy (thecalchater):

ok let me do that real quick

OpenStudy (welshfella):

what do you get for that?

OpenStudy (thecalchater):

that is false... You didn't list a variable and 15000 doesn't equal the opposite side.

OpenStudy (welshfella):

what ?? the left side is the required mass

OpenStudy (welshfella):

thats the variable

OpenStudy (thecalchater):

15,000 = 11 * e^(-0.00043349* 15,000) this is what u put did u mean to put an x where the 15000 is ?

OpenStudy (welshfella):

work out the right side.

OpenStudy (thecalchater):

Then I got A or .0165

OpenStudy (welshfella):

thats it

OpenStudy (thecalchater):

ok moving on to 2 and 3? Which one do u want to do first?

OpenStudy (welshfella):

i gotta go right now . I#m sure you#ll get help from somebody

OpenStudy (thecalchater):

ok thanks for the help! You guys that are here do u want em to open another question with only i listed or do u want to keep it going here?